OpenStudy (anonymous):

Solve each system graphicaly. be sure to check your solution. if a system has an infinite number of solutions, use set-builder notation to write a solution set. If a system has no solution, state this. 4x - y = 9, x - 3y =16

OpenStudy (saifoo.khan):

i got the solutions! :O

OpenStudy (anonymous):

answer or method?

OpenStudy (anonymous):

method

OpenStudy (saifoo.khan):

lol, i got both!

OpenStudy (anonymous):

multiply the first equation all the way across by -3

OpenStudy (anonymous):

\[-12x+3y=-27\]

OpenStudy (anonymous):

then add in a line to the two equations together. you get \[-12x+3y=-27\] \[x-3y=16\] \[-11x=-11\] \[x=1\]

OpenStudy (anonymous):

now that you know x = 1 replace x by 1 in either equation to find y

OpenStudy (anonymous):

the first one gives \[4\times 1-y=9\] \[4-y=9\] \[-y=5\] \[y=-5\]

OpenStudy (anonymous):

so solution is \[x=1,y=-5\] or \[(1,-5)\]